# Let the Trial Begin: A Mock-Court Approach to Vulnerability Detection using LLM-Based Agents

### v1

@@ 4.3. RQ3: Effect of Multiple Rounds of Discussion
For GPT-4o, increasing the number of rounds actually led to a decrease in overall performance. Specifically, the proportion of pairs incorrectly categorized as vulnerable (P-V) increases with more rounds. From further analysis, we found that the security researcher agent suggested fewer potential vulnerabilities in later rounds, with an average of 4.22 per code in Round 1, decreasing to 3.55 in Round 2 and 3.43 in Round 3. This decline is also driven by the code author agent’s refutations. In 30.88% out of the total prediction, the security researcher removed a previously identified vulnerability after the code author disputed it. Meanwhile, the code author agent became progressively more inclined to acknowledge vulnerabilities, with its mitigation-to-refutation ratio shifting from 1.73:1 in Round 1, to 2.5:1 in Round 2, and finally 3:1 in Round 3. Although the review board agent did not blindly accept arguments from either side (for instance, during the three-round setup, in 100 out of 868 code instances, review board agent overruled the code author’s refutation and decided the vulnerability to be valid), the overall imbalance introduced by these interactions still raised P-V and lower P-C rates.
@@
Answer to RQ3 Using GPT-4o in VulTrial outperforms both single-agent and multi-agent baselines across different rounds of agent interaction (i.e., 1, 2, and 3). However, performance declines when extending beyond a single discussion round.
@@
Base 22.07 35.40 30.80 11.72
No moderator 7.36 8.51 77.70 6.44
No code author 14.48 63.68 11.95 9.89
No moderator+code author 14.48 61.84 14.48 9.20
No moderator+review boards 2.07 93.56 1.15 3.22
4.4. RQ4: Ablation Study
In this RQ, we investigate the necessity of each agent in VulTrial. Table 4 presents the ablation results under different agent-removal scenarios. Overall, we find that all four agents are essential for achieving the best performance.
Removing the moderator agent from VulTrial reduces P-C from 22.07 to 7.36. This result highlights the moderator’s critical role in summarizing arguments to help review board agent in making the decision. This finding is also consistent with the results of RQ2 (see Section 4.2), where instruction tuning the moderator yielded the largest performance gain. Removing both the moderator and review board agent causes P-C to further drop 2.07, while P-V (pairs both labeled vulnerable) increase to 93.56. Highlighting that most of the codes are predicted as vulnerable by the code author. This imbalance occurs because, without an additional agent review board, the final judgment defaults heavily toward labeling codes as vulnerable based on partial agreements from the code author agent. Thus, the review board is essential for ensuring balanced final verdicts.
Removing the code author agent from VulTrial reduces P-C to 14.48 and increases P-V to 63.68. Without the code author’s counterarguments, the review board agent has a limited perspective and frequently accepts vulnerabilities identified by the security researcher agent. Removing both the moderator and code author produces similar results (P-C: 14.48, P-V: 61.84). This underscores the crucial role of the code author in ensuring balanced judgments while also demonstrating that the moderator agent becomes redundant without opposing arguments to mediate.
Answer to RQ4 Our findings confirm that every agent in VulTrial is necessary to achieve the best performance.
@@ 5. Discussion
By assigning specialized roles to agents, mirroring real-world scenarios, VulTrial can facilitate deeper discussions by incorporating diverse rationales from different parties, even when using small-scale LLMs. However, this multi-agent approach inherently incurs additional costs due to the increased number of input and output tokens. Interestingly, while GPT-4o is generally more expensive and performs better than GPT-3.5 as a single-agent, with respective P-C 9.2 and 6.21, using a multi-agent approach (VulTrial) with GPT-3.5 outperforms GPT-4o’s single-agent performance by 69.89%. This approach also reduces costs, requiring only \$7.46 compared to GPT-4o’s single-agent cost of \$8.10 for the whole testing data. Conversely, applying VulTrial to more expensive models, such as GPT-4o, significantly increases costs (approximately \$23.9=32-8.1). Therefore, careful consideration is important when selecting models to optimize both cost efficiency and detection performance.
Multi-Agents Performance
In both experiments using GPT-3.5 and GPT-4o as base models for GPTLens (see Section 4.1), we observed that a large portion of the codes were predicted as vulnerable, as shown by high P-V scores (incorrectly predicted as vulnerable) of 89.2 and 41.15, respectively. This behavior may be due to the critic agent (final agent) assigning high scores on potential vulnerabilities flagged by the auditor agent (first agent). To get the final prediction, GPTLens also utilizes a threshold to determine whether a function is vulnerable. The final score, ranging from 0 to 9, is computed as $0.5\times\text{correctness}+0.25\times\text{severity}+0.25\times\text{profitability}$, with correctness, severity, and profitability taken from the critic agent output.
The final scores below 5 are considered benign. When GPTLens uses GPT-3.5, the median final score is 7.25 with mean 7.08, indicating that most code is labeled as vulnerable. Similarly, when GPTLens uses GPT-4o, the median final score is 6.75 with mean 6.67, which means the majority of the code prediction also exceeds the benign threshold.
Meanwhile, under GPT-4o, VulTrial demonstrated a more balanced results, as indicated by a mitigation-to-refutation ratio of 2:1 in the code author agent’s responses. Furthermore, we found that in 101 benign and 253 vulnerable codes, the code author agent refuted at least one potential vulnerability initially flagged by the security researcher (i.e., disagreement between code author and security researcher), yet the final label remained correct. These cases, accounting for 40.78% of the total codes, demonstrate VulTrial’s effectiveness in resolving disagreements.
Enhancing Role-Specific Agent
@@
Tuning the moderator agent in VulTrial using GPT-4o enhances single-agent performance by 139.89% and improves the non-tuned version of VulTrial by 18.52%. This underscores the importance of instruction tuning for role-specific agents. While instruction tuning on all four agents in VulTrial would cost four times more than tuning a single-agent. In this study, we used only a sample of 50 pairs for instruction tuning, which kept the cost at approximately \$17 for each agent. This is significantly lower than using the instruction tuning setting done in PrimeVul (Ding et al., 2025b), which would cost approximately more than \$500 for GPT-4o (¿15,000 samples).
@@ 6. Threats to Validity
External Validity. An external threat to our study is the generalizability of our findings. In this study, we focused the evaluation using PrimeVul pair data, which is a challenging recently-released vulnerability detection task on C/C++ programming language. We did not use other datasets as they may not be free from bias that have been uncovered by Ding et al. (Ding et al., 2025b). In the future, we plan to investigate VulTrial performance using other clean datasets.
Construct Validity. Threats to construct validity are from the choice of metrics in our study. In this study, we utilize P-C, P-B, P-V, P-R (Section 3.4). These metrics are designed for a pair-wise evaluation introduced in PrimeVul (Ding et al., 2025b). In a textually similar pair of functions, the metrics assesses an approach’s ability to correctly identify both the presence and absence of vulnerabilities. Consequently, we believe that the threat to construct validity is minimal.
Internal Validity. Threats to internal validity in our study refer to possible errors in our experiment. To mitigate the risk, we have carefully checked the experiments that we run in our study. The code, prompt, and results are also included in our replication package: https://figshare.com/s/1514bc9a7aa64b46d94e. Furthermore, to reduce the randomness effect of the experiment, we set the temperature to 0 to make the results more deterministic. We also did the preliminary evaluation where we ran the VulTrial on testing data three times and found that the standard deviation is low (0.7).
@@ 7. Related Work
Single-Agent Vulnerability Detection. A broad range of studies have explored prompting and CoT for vulnerability detection (Tamberg and Bahsi, 2024; Ullah et al., 2023; Zhou et al., 2024; Nong et al., 2024; Zhang et al., 2024; Ding et al., 2025b; Steenhoek et al., 2024). Tamberg and Bahsi (Tamberg and Bahsi, 2024) benchmarked LLMs using advanced prompting techniques, including CoT, tree-of-thought, and self-consistency, finding that CoT with GPT-4 achieved the best performance and outperformed static analysis tools. Similarly, Zhou et al. (Zhou et al., 2024) demonstrated that GPT-4 with prompting outperforms fine-tuned CodeBERT. Ullah et al. (Ullah et al., 2023) evaluated commercial LLMs across various prompting strategies, including CoT, and found promising detection performance. However, they noted often unfaithful reasoning in LLMs leads to incorrect classifications. Recent studies have further highlighted the limitations of CoT (Nong et al., 2024; Zhang et al., 2024; Steenhoek et al., 2024). Notably, PrimeVul’s seminal work found that GPT-3.5 and GPT-4 perform poorly on PrimeVul’s pair test set, reinforcing concerns about CoT’s reliability in vulnerability detection. In this work, we use CoT as a baseline and show that our approach, VulTrial, significantly outperforms it across various LLMs on PrimeVul.
@@
Multi-Agent Vulnerability Detection. Multi-agent frameworks can enhance LLM-based vulnerability detection by enabling collaboration and specialization among agents (He et al., 2024). For example, EvalSVA (Wen et al., 2024) introduces a multi-agent framework where agents act as security experts that improves over single-agent LLMs. Similarly, MuCoLD (Mao et al., 2024) adopts a multi-agent setup where agents assume roles such as tester and developer, which collaboratively analyze code. In contrast, VulTrial employs a structured multi-agent design inspired by a courtroom, where agents take on distinct roles such as prosecutor, defender, and judge. This courtroom analogy encourages adversarial debate and balanced reasoning that leads to robust decision-making. We excluded these recent works as baselines because the authors did not provide a replication package for reusing their agents, and the prompt information in the paper was insufficient to replicate the study.
@@
Other multi-agent-based approaches have been explored for smart contract auditing (Wei et al., 2024; Ma et al., 2024; Hu et al., 2023). LLM-SmartAudit (Wei et al., 2024) uses collaborative decision-making with role-play and reasoning to detect vulnerabilities in smart contracts. Similarly, iAudit (Ma et al., 2024) assigns specific roles, such as detector, reasoner, and critic, to guide the analysis of Solidity smart contracts. GPTLens (Hu et al., 2023) employs multiple LLM-based agents acting as independent prosecutors to identify vulnerabilities, with a critic agent reviewing and ranking the findings. Our work differs from these studies in two key ways. First, we focus on vulnerability detection in C and C++ on PrimeVul, a benchmark comprising challenging real-world examples. Second, VulTrial adopts a structured, courtroom-inspired framework specifically tailored for general-purpose languages beyond smart contracts. As these studies focus on smart contracts, their scope differ from our work for general-purpose programming languages. Nonetheless, we found that GPTLens can be adapted to general-purpose vulnerability detection and included it as a baseline in our experiments (Section 3.3). Our results demonstrate that VulTrial substantially outperforms GPTLens and single-agent baselines on PrimeVul.
8. Conclusion and Future Direction

### v2

@@ 4.3. RQ3: Effect of Multi Rounds Discussion
In the previous RQs, VulTrial used a single round of interaction among the agents. In this RQ, we investigate whether additional discussion rounds enhance performance. Table 3 shows the results of these multiple rounds experiments, showing that the best result is achieved using a single round interaction. For GPT-4o, increasing the number of rounds actually led to a decrease in overall performance. Specifically, P-V increases with more rounds. From further analysis, we found that the security researcher agent suggested fewer potential vulnerabilities in later rounds, with an average of 4.22 per code in Round 1, decreasing to 3.55 in Round 2 and 3.43 in Round 3. This decline is also driven by the code author agent’s refutations. In 30.88% of the total predictions, the security researcher removed a previously identified vulnerability after the code author disputed it. Meanwhile, the code author agent became progressively more inclined to acknowledge vulnerabilities, with its mitigation-to-refutation ratio shifting from 1.73:1 in Round 1, to 2.5:1 in Round 2, and finally 3:1 in Round 3. Although the review board agent did not blindly accept arguments from either side (for instance, during the three-round setup, in 100 out of 868 code instances, review board agent overruled the code author’s refutation and decided the vulnerability to be valid), the overall imbalance introduced by these interactions still raised P-V and lower P-C rates.
@@
Answer to RQ3: Using GPT-4o in VulTrial outperforms both single-agent and multi-agent baselines across different rounds of agent interaction (i.e., 1, 2, and 3). However, performance declines when extending beyond a single discussion round.
4.4. RQ4: Ablation Study
In this RQ, we investigate the necessity of each agent in VulTrial. Table 4 presents the ablation results under different agent-removal scenarios. Overall, we find that all four agents are essential for achieving the best performance. Removing the moderator agent from VulTrial reduces P-C from 96 to 32. This result highlights the moderator’s critical role in summarizing arguments to help review board agent in making the decision. This finding is also consistent with the results of RQ2 (see Section 4.2), where instruction tuning the moderator yielded the largest performance gain. Removing both the moderator and review board agent causes P-C to further drop to 9, while P-V (pairs both labeled vulnerable) increases to 407. Highlighting that most of the codes are predicted as vulnerable by the code author. This imbalance occurs because, without an additional agent review board, the final judgment defaults heavily toward labeling codes as vulnerable based on partial agreements from the code author agent. Thus, the review board is essential for ensuring balanced final verdicts.
Removing the code author agent from VulTrial reduces P-C to 63 and increases P-V to 277. Without the code author’s counterarguments, the review board agent has a limited perspective and frequently accepts vulnerabilities identified by the security researcher agent. Removing both the moderator and code author produces similar results (P-C=63, P-V=269). This underscores the crucial role of the code author in ensuring balanced judgments while also demonstrating that the moderator agent becomes redundant without opposing arguments to mediate.
Answer to RQ4: Our findings confirm that every agent in VulTrial is necessary to achieve the best performance.
@@ 4.4. RQ4: Ablation Study
Base 96 154 134 51
No moderator 32 37 338 28
No code author 63 277 52 43
No moderator+code author 63 269 63 40
No moderator+review boards 9 407 5 14
@@ 5. Discussion
By assigning specialized roles to agents, mirroring real-world scenarios, VulTrial can facilitate deeper discussions by incorporating diverse rationales from different parties. However, this multi-agent approach inherently incurs additional costs due to the increased number of input and output tokens. Interestingly, while GPT-4o is generally more expensive and performs better than GPT-3.5 as a single-agent, using VulTrial with GPT-3.5 outperforms GPT-4o’s single-agent P-C by 28. This approach also reduces costs, requiring only \$7.46 compared to GPT-4o’s single-agent cost of \$8.10 for the whole testing data. Conversely, applying VulTrial to more expensive models, such as GPT-4o, significantly increases costs (approximately \$23.9=32-8.1). Therefore, careful consideration is important when selecting models to optimize both cost efficiency and detection performance.
Enhancing Role-Specific Agents
@@
Tuning the moderator agent in VulTrial using GPT-4o improves the baselines and non-tuned version of VulTrial. These results highlight the potential for improving performance by tuning role-specific agents. Our study used only a sample of 50 pairs for instruction tuning, keeping the cost at approximately \$17 for each agent. This is significantly lower than using the instruction tuning setting done in PrimeVul (ding2024vulnerability), which would cost more than \$500 for GPT-4o (¿15,000 samples).
Analysis on Instruction Tuning Data.
Comparing the 50 fine-tuning pair samples from the validation dataset with the 96 successful VulTrial pairs, we found that 39 projects from the successful cases do not appear in the fine-tuning set. Only 8 projects are shared. Of these, 3 projects contain only commits that occurred after the corresponding training commits. In total, only 11 successful pairs involve commits made earlier than the latest commit used in training. The fine-tuning data includes 24 unique CWEs, 20 of which also appear in the successful set. 12 CWEs in the successful set are not covered by the fine-tuning data. These results show minimal overlap in terms of projects, commit timing, and CWE types, suggesting that the model’s performance is not a result of memorization.
Table 5. Results of running VulTrial in the wild.
Project Funcs analyzed Alarms raised False positives True positives
Superset (superset) 47 0 0 0
DataHub (datahub) 160 2 0 2
InvenTree (inventree) 11 1 0 1
VulTrial in the Wild
To assess VulTrial in real-world settings, we conducted a pilot experiment on selected open-source projects. Specifically, we targeted well-maintained repositories with active security practices. To identify such projects, we filtered GitHub repositories that include a SECURITY.md file and have over 1,000 stars. We chose the projects that have a SECURITY.md file, as it typically indicates that maintainers are attentive to security concerns and have set up a process for reporting vulnerabilities.
This selection process yielded 196 repositories. From this set, we randomly chose three projects for analysis. For each project, we analyzed recent commits, extracted modified functions, and applied VulTrial with its default rule set, no project-specific tuning. The results of this experiment are summarized in Table 5. We manually inspected the flagged vulnerabilities, attempted exploit generation, and reported confirmed issues to the respective maintainers if the vulnerable code was used in the project (i.e., not dead code).
In InvenTree, we discovered a denial-of-service vulnerability in the built-in label_sheet plugin. The plugin accepts a skip parameter with no upper bound, allowing an attacker to allocate excessively large Python lists and exhaust server memory. This issue affects any authenticated user who can request labels and has been assigned CVE-2025-49000. In DataHub, VulTrial identified two functions with resource-exhaustion risks (CWE-400). One function is still actively used, while the other is no longer in use but remains in the codebase. We disclosed the first issues following instructions on their SECURITY.md, and the DataHub developers confirmed the findings. These results demonstrate that VulTrial is capable of detecting previously unknown vulnerabilities in the wild. To reduce the burden on maintainers, we adopted a conservative reporting strategy and only reported issues we could confidently confirm as true vulnerabilities. We have made the results, vulnerability reports, and exploits available (ourreplicationpackage).
Generalizability of VulTrial
In our main experiments, VulTrial, using OpenAI’s models, outperforms the baseline methods. Next, to investigate its generalizability, we further evaluated it with a different model family, specifically, LLaMA-3.1-8B (dubey2024llama). Table 6 shows the results, indicating the consistent performance of VulTrial even with a different model family. CoT struggles under LLaMA-3.1-8B, with only 4 correct predictions (P-C) and a recall of 0.02, largely due to a high number of benign misclassifications (P-B = 425). VulTrial improves recall to 0.48 and raises P-C to 89, although it has an increased false-positive rate (FPR = 0.45). This suggests that VulTrial favors broader vulnerability coverage, potentially at the cost of mislabeling some benign pairs.
When compared with GPTLens, VulTrial presents a different trade-off. GPTLens achieves higher recall (0.55) and fewer benign misclassifications (P-B = 103), but it also produces more P-V (158) and a higher FPR (0.58). VulTrial, in contrast, has more correct predictions overall (P-C = 89 vs. 80) and higher precision (0.52 vs. 0.49), while maintaining a lower false-positive rate (0.45). These differences indicate that VulTrial may provide a more conservative alternative to GPTLens, with improved balance across evaluation metrics. Overall, these findings suggest that VulTrial’s multi-agent reasoning approach can generalize to other models.
Table 6. Results of VulTrial, single-agent, and multi-agent baselines with LLaMA-3.1-8B as the base model.
[…]
CoT(ding2024vulnerability) 4 3 425 3 431 0.54 0.02 0.01
GPTLens(gptlens) 80 158 103 94 355 0.49 0.55 0.58
VulTrial 89 120 150 76 346 0.52 0.48 0.45
Evaluating Explanation
VulTrial provides explanations for each agent’s decisions, helping developers understand why a function is classified as vulnerable or benign rather than receiving a simple binary outcome. Because multiple agents contribute, developers may need to read the full multi-agent discussion to grasp the reasoning. To improve readability, we use GPT-4o to synthesize all agent reasoning into a single coherent explanation. We compare the synthesized VulTrial explanations with CoT-based ones using four metrics (i.e., completeness, informativeness, clarity, and actionability), each scored on a 1-5 scale (widyasari2025explaining; degen2022explain; zhou2021evaluating; velmurugan2025developing). For fairness, we evaluate only cases where the P-C intersections of VulTrial and CoT match under GPT-4o. We conduct automated evaluation using a closed-source LLM from a different family (Claude-Haiku-4.5 (anthropic2025claudeHaiku)) and manual evaluation by three external annotators on 20 instances, without revealing explanation sources. Both evaluations show that synthesized VulTrial explanations outperform CoT across all metrics (Table 7). All generated explanations and evaluation results are available in the replication package (ourreplicationpackageall).
@@ 6. Threats to Validity
External Validity. In this study, we focused the evaluation using PrimeVul pair data, which is a challenging recently-released vulnerability detection task on C/C++. We did not use other datasets as they may not be free from bias that have been uncovered by Ding et al. (ding2024vulnerability). In the future, we plan to investigate VulTrial performance using other clean datasets. Another threat is the risk of data leakage, as ChatGPT’s training data is not publicly disclosed. We mitigate this by evaluating VulTrial using LLaMA-3.1-8B, whose training cutoff (Dec 2023) is before PrimeVul’s GitHub release (Mar 2024). VulTrial also identifies previously unknown real-world vulnerabilities, including CVE-2025-49000, showing it can generalize beyond training data. Therefore, we believe the threats are minimal.
Construct Validity. In this study, we use both pair-wise metrics and traditional metrics in our experiments (Section 3.4). These metrics provide a reliable measure of comparative performance, minimizing the threat to construct validity. In our experiments, VulTrial uses a single underlying model for all agents (e.g., GPT-4o, and LLaMA-3.1-8B), we rely on prompt-based role specialization to induce diverse behaviors (e.g., prosecutor vs. defense attorney). Prior works (wei2022chain; chatterjee2024posix; hu2023large; shinn2023reflexion; wu2024autogen) have shown that such prompt-induced roles can lead to qualitatively different reasoning paths even within the same model. In future work, we plan to use different LLMs for each agent role to further enhance diversity.
Internal Validity. Threats to internal validity in our study refer to possible errors in our experiment. To mitigate the risk, we have carefully checked the experiments that we run in our study. The code, prompt, and results are also included in our replication package (ourreplicationpackageall). Furthermore, we set the temperature to 0 to reduce the randomness effect of the experiment. We also did the preliminary evaluation where we ran the VulTrial on testing data three times and found that the standard deviation is low (0.7).
Table 7. Results of automatic and manual evaluations of the generated explanation from CoT and synthesized VulTrial.
CoT / Synthesized VulTrial
Completeness Clarity Actionability Inform.
Claude (anthropic2025claudeHaiku) 3.15 / 4.10 3.60 / 3.85 2.15 / 4.00 2.80 / 3.95
Human 3.83 / 4.32 3.72 / 4.20 2.37 / 4.52 3.12 / 3.63
@@ 7. Related Work
Single-Agent Vulnerability Detection. Recent work has explored both DL-based and LLM-based methods for vulnerability detection. DL approaches, such as LineVul (fu2022linevul), achieve high recall with the use of transformer-based line-level classification. Ni et al. (ni2023distinguishing) proposed a semantic representation learning method, while Steenhoek et al. (steenhoek2024dataflow) incorporated dataflow. In contrast, LLM-based methods have gained traction for their strong results (gptlens; zhou2024large_emerging). Other studies have explored prompting and CoT (tamberg2024harnessing; ullah2023llms; zhou2024large_emerging; nong2024chain; zhang2024prompt; ding2024vulnerability; steenhoek2024comprehensive). Tamberg and Bahsi (tamberg2024harnessing) benchmarked LLMs using advanced prompting techniques, including CoT, tree-of-thought, and self-consistency, finding that CoT with GPT-4 achieved the best performance and outperformed static analysis tools. Similarly, Zhou et al. (zhou2024large_emerging) demonstrated that GPT-4 with prompting outperforms fine-tuned CodeBERT. Ullah et al. (ullah2023llms) evaluated commercial LLMs across various prompting strategies, including CoT, and found promising detection performance. Recent studies have further highlighted the limitations of CoT (nong2024chain; zhang2024prompt; steenhoek2024comprehensive). Notably, PrimeVul’s seminal work found that GPT-3.5 and GPT-4 perform poorly on PrimeVul’s pair test set, reinforcing concerns about CoT’s reliability in vulnerability detection. Efforts to enhance LLM-based methods include finetuning and retrieval-augmented generation (RAG). Shestov et al. (shestov2025finetuning) showed that finetuning boosts recall but is costly and dataset-sensitive. Du et al. (du2024vul) proposed Vul-RAG, which improves factual grounding via structured knowledge injection. Sun et al. (sun2024llm4vuln) introduced LLM4Vuln, a framework that evaluates and decouples LLM reasoning capabilities. In this work, we use DL-based methods and CoT as baselines and show that VulTrial outperforms them on PrimeVul.
@@
Multi-Agent Vulnerability Detection. Multi-agent frameworks can enhance LLM-based vulnerability detection by enabling collaboration and specialization among agents (he2024llm). EvalSVA (wen2024evalsva) introduces a multi-agent framework where agents act as security experts. Similarly, MuCoLD (mao2024multi) agents assume roles such as tester and developer, which collaboratively analyze code. In contrast, VulTrial employs a structured multi-agent design inspired by a courtroom, where agents take on distinct roles such as prosecutor, defender, and judge. This encourages adversarial debate and balanced reasoning that leads to robust decision-making. We excluded these recent works as baselines due to the lack of replication packages and insufficient prompt details to replicate the study.
@@
Other multi-agent-based approaches have been explored for smart contract auditing (wei2024llm; ma2024combining; hu2023large). LLM-SmartAudit (wei2024llm) uses collaborative decision-making with role-play and reasoning to detect vulnerabilities in smart contracts. Similarly, iAudit (ma2024combining) assigns specific roles, such as detector, reasoner, and critic, to guide the analysis of Solidity smart contracts. GPTLens (hu2023large) employs multiple LLM-based agents acting as independent prosecutors to identify vulnerabilities, with a critic agent reviewing and ranking the findings. Our work differs from these studies in two key ways. First, we focus on vulnerability detection in C and C++ on PrimeVul, a benchmark comprising challenging real-world examples. Second, VulTrial adopts a structured, courtroom-inspired framework specifically tailored for general-purpose languages beyond smart contracts. As these studies focus on smart contracts, their scope differs from our work for general-purpose programming languages. Nonetheless, we found that GPTLens can be adapted to general-purpose vulnerability detection and included it as a baseline in our experiments (Section 3.3). Our results demonstrate that VulTrial substantially outperforms GPTLens and single-agent baselines on PrimeVul.
8. Conclusion and Future Work